(c20p24) A single loop of area 0.193 m2 is placed perpendicular to a magnetic field of 0.334 T. The loop is rotated about an axis lying in the plane of the loop and passing through its center. An instantaneous emf of 8.21 Volts is generated when the plane of the loop makes an angle of 45 o with the field direction. What is the angular speed of the loop (rad/s)?
Answer #1

Faraday's law:

V = N*d(BA)/dt

= NB*dA/dt

= NBA*w*sin(theta)

Thus angular rate

w = V/(NBA*cos(theta))

w = 9.81/(1*0.334*0.193*cos(45 deg)) = 215.21 rad/s